By referring to the chemical equation given, calculate the volume of nitrogen gas in litres, required to react with 40.0 g of magnesium at a pressure of 2.25 atm and a temperature of 15 °C.

            3Mg (s) + N2 → Mg3N2 (s)


1
Expert's answer
2021-10-16T02:02:14-0400

Use the ideal gas equation, PV = nRT find the volume of nitrogen gas in liters from moles of N2.


Step 1: Convert 40.0 g of Mg to moles by using the atomic mass of Mg.


Atomic mass of Mg = 24.305 g/mol




"moles \ of \ Mg \ = \ 40.0 \cancel{\ g \ of \ Mg} \ * \ \frac{1 \ mol \ of \ Mg}{24.305 \ \cancel{g \ of \ Mg} }""= \ 1.6458 \ mol \ of \ Mg."

Step 2: Using the mol to mol ratio of Mg and N2 from the balanced chemical reaction find the moles of N2 that will react with 1.6458 moles of N2.


  3Mg (s) + N2 → Mg3N2 (s)


The mol to mol ratio of Mg and N2 in this balanced chemical reaction is 3: 1.





"moles \ of \ Mg \ = \ 1.6458 \cancel{\ mol \ of \ Mg} \ * \ \frac{1 \ mol \ of \ N_2}{3 \ \cancel{mol \ of \ Mg} }""= \ 0.5486 \ mol \ of \ N_2"

Step 3: Using the ideal gas equation, PV = nRT find the volume of nitrogen gas in liters occupied by 0.5486 moles of N2 at a pressure of 2.25 atm and a temperature of 15 °C.


Pressure, P = 2.25 atm,


Temperature, T in Kelvin = 15 °C = 15 + 273.15 = 288.15 K


moles of N2 , n = 0.5486 moles of N2


Gas constant, R = 0.08206 L.atm/mol.K


plug all this information in the ideal gas equation we have


"2.25 \ atm * V \ = \ 0.5486 \ mol \ * \ 0.08206 \ L.atm/mol.K * 288.15 \ K"


"2.25 \ atm * V \ = \ 12.9716 \ L.atm"


dividing both the side by 2.25 atm, we have




"\frac{\cancel{2.25 \ atm} * V}{ \cancel{2.25 \ atm}} = \ \frac{12.9716 \ L.atm}{2.25 \ atm }"




"V = 5.765 \ L"